# Transfer tensor analysis of localization in the Anderson and Aubry-André-Harper models

### v1

@@ Abstract
We use the transfer tensor method to analyze localization and transport in simple disordered systems, specifically the Anderson and Aubry-André-Harper models. Emphasis is placed on the memory effects that emerge when ensemble-averaging over disorder, even when individual trajectories are strictly Markovian. We find that memory effects are necessary to remove fictitious terms that would correspond to redrawing static disorder at each time step, which would create a temporally uncorrelated dynamic disorder. Our results show that while eternal memory is a necessary condition for localization, it is not sufficient. We determine that signatures of localization and transport can be found within the transfer tensors themselves by defining a metric called “outgoing-pseudoflux”. This work establishes connections between theoretical research on dynamical maps and Markovianity and localization phenomena in physically realizable model systems.
@@ I Introduction
Theoretical study of quantum transport in realistic systems is complicated by the unfavorable scaling of computational requirements when modeling quantum systems,[14] so simple but powerful treatments such as the Anderson,[11], Haken-Strobl-Reineker (HSR),[15, 16] or Aubry-André-Harper (AAH) models[17, 18] are often employed to understand transport and localization in tight binding lattices. As many experimental measurements sample over regions with defects, inducing effects such as inhomogenous broadening on spectroscopic measurements,[19, 1, 20, 21, 22] it is common practice to average over static disorder in model parameters in order to better describe experimental observations.[11, 23, 24, 25] These simple models have increased our understanding of phenomena including metal-insulator transitions and diffusion rates in the presence of thermal noise and disorder.[26, 27, 28, 23, 29, 17, 18, 24]
The dynamics resulting from averaging over a static disorder ensemble recovers the spatial homogeneity that was destroyed when the disorder was introduced. However, a somewhat counterintuitive complication arises. Evolutions consisting of mixtures of distinct unitary evolutions, averages over static disorder in our case, may show memory effects. Although the exact interpretation of these effects, and of quantum non-Markovianity in general, remains a matter of discussion,[30, 31, 32, 33, 34] the association of this apparent non-Markovianity and Anderson localization has been noted in previous studies.[35, 36, 37] We analyze the origin of these memory effects using transfer tensors (TT), which are ideally adapted to this purpose. We see that memory terms differentiate static disorder from dynamic disorder and find that, although an infinite memory time does not indicate whether a model undergoes localization, another characteristic of the transfer tensors, which we call outgoing-pseudoflux, provides an indication of localization, demonstrating the potential of TT as a diagnostic tool for assessing transport properties and understanding the origin and impact of memory. This method further allows us to illustrate the links between physically realizable systems such as the AAH model[38, 39, 40, 41, 42, 43] and more abstract discussions about quantum non-Markovianity and convex combinations of dynamical maps.[30, 31, 32, 33, 34]
@@ II Transfer Tensor Analysis of the Anderson Model
in the lattice site basis, $|n\rangle$, which includes nearest-neighbor interactions with coupling strength $V$ and site energies, $\epsilon_{n}$, drawn from a normal distribution with zero mean and standard deviation $\sigma$. Throughout this paper, we set $\hbar=1$ and $V=1$. Introducing static disorder localizes the system in 1D and breaks spatial homogeneity,[11, 44, 1] but averaging over many realizations of the static disorder restores it. Whereas the individual trajectories in the ensemble are Markovian and have no memory, the spatially homogeneous ensemble evolution now has memory, as noted in the context of convex mixing of quantum channels or dynamical maps[31, 14] and in previous studies of Anderson localization.[35, 36] To better see this, we adopt the transfer tensor method to describe the disorder-averaged dynamics.
@@
We will find that individual elements of the transfer tensor can provide useful information. Because we average over disorder and restore spatial homogeneity, we can analyze only the transfer tensor elements involving the center site of the system. The element $T_{m,n,o,p}(l)$ quantifies the influence that $\rho_{o,p}(k)$ will have on $\rho_{m,n}(k+l)$ and $T_{m,n,o,p}(l)$ depends only on the differences mpm-p, npn-p, and opo-p. The utility of an equivalent relationship for the generalized quantum master equations was recently recognized and exploited for performance enhancement. [47] The spatial homogeneity of the TT allows analysis with very modest basis sizes.
@@
Figure 1: All calculations in this figure involved a basis of size 1515, $\Delta t=V/10$ and 20002000 trajectories. (a) The Frobenius norm of the transfer tensor $T(k)||$ as a function of $k$ for several $\Gamma=0$ with different $\sigma$ strengths. Note that the norm of the first transfer tensor is on the order of the basis size and not shown on these axes. (b) Population of site $m=8$ following initialization in a delta function localized on $m$ as a function of time step for several ensembles including a case in which all transfer tensor memory terms were artificially set to zero. The dashed black line is the equilibrium population. (c) The Frobenius norm of the transfer tensor $T(k)||$ as a function of time step $k$ for several different $\Gamma$ strengths with $\sigma=V$. Note that the norm of the first transfer tensor is on the order of the basis size and not shown on these axes. (d) The cumulative outgoing-pseudoflux as a function of time step for the same $\Gamma$ strengths and $\sigma$ as in (c) showing convergence to non-zero values at finite $\Gamma$ and convergence to zero at zero $\Gamma$. Note that, when $\Gamma=10$, noise overpowers disorder and the system is effectively memoryless for the chosen $\Delta t$.
The memory that arises for the disorder-averaged ensembles is quantified in Fig. 1 (a) which shows the Frobenius norms of the TT, $T(k)||$, as a function of $k$ for several $\sigma$. In the case where $\sigma=0$, memory is zero within the numerical tolerance for all $k>1$. For $\sigma>0$, the tensor norms converge toward a finite value as $k\to\infty$. This finite value increases with $\sigma$ and suggests an eternal memory. This is analogous to the eternally non-Markovian quantum master equation seen in the literature.[34] There is no information backflow from an environment to the system in this model as there is no environment at all. This indicates that the disordered Anderson evolution is, by some definitions of quantum non-Markovianity, Markovian despite the necessity of describing memory effects.[48, 33, 34]
@@
Upon closer examination, the Markovian evolution represented by $M(k)=[T(1)]^{k}$ in the above discussion is equivalent to the dynamics induced by averaging uncorrelated Gaussian white noise in the limit of $\Delta t\rightarrow 0$. This is actually the HSR model, averaged over stochastic noise,[16, 15] where the site energies follow zero mean Gaussian statistics
@@
This gives rise to the equivalent Gorini–Kossakowski–Sudarshan–Lindblad equation of motion for the corresponding noise-averaged density matrix,
@@
When the HSR model is combined with the disordered Anderson Hamiltonian the dynamics shows diffusive transport whenever $\Gamma$ is nonzero.[24, 49, 50, 51] We examine the interplay between static disorder and dynamic noise with respect to transport further using the TT formalism. One can readily deduce that in the limit of an infinitesimal time step, $M(1)$ will be identical for all ensembles for which $\Gamma+\sigma^{2}\Delta t=C$, with $C$ an arbitrary constant. Fig. 1 (b) shows the population as a function of time at the center site for a set of ensembles with $C=0.1$, in which all dynamics are initialized as a Kronecker delta. In one special case, an ensemble was prepared for $\sigma=1$ and $\Gamma=0$, but the resulting memory terms were artificially set to zero, i.e. $T(k)=0$ for all $k>1$ while keeping the first term the same. The resulting dynamics is identical to a $\sigma=0$, disorder-free and naturally memoryless propagation with $\Gamma=0.1$. This demonstrates that when memory terms are eliminated, static and dynamic disorder cannot be differentiated. In Fig. 1 (b), several time steps are required before memory effects begin to manifest impacts on M(i)M{(i)} and the populations from different ensembles diverge, with divergence occurring more quickly for larger $\sigma$. Regardless of the size of $\sigma$, as long as $\Gamma$ is nonzero all populations will eventually converge to the infinite temperature equilibrium state where all sites are equally occupied, indicated with a dashed black line in the figure.
@@
Second, inspection of individual tensor elements can indicate transport properties. It is customary to follow the time-dependent spatial profile with certain localized initial conditions as in microscopy experiments to assess transport.[52, 53] To facilitate comparisons between different systems, averaged measures such as mean squared displacement $\langle x^{2}\rangle$, whose scaling with time can be used to classify the nature of transport, e.g. $\langle x^{2}\rangle\propto t$ for diffusive transport, are employed. Here, we take a different approach and define the outgoing-pseudoflux at time step $k$ for site $m$ as $f_{m}(k)=\sum_{n}^{\prime}{T}_{n,n,m,m}(k),\;\;n\neq m$ and its cumulative sum over time $F_{m}(k)=\sum_{l=1}^{k}f_{m}(l)$. The former represents the magnitude of the influence that the population of the $m$th site $k$ steps in the past has on the populations of all other sites in the system at the current time step. The physical significance of $F_{m}(k)$ is further discussed in Appendix A, where we illustrate the association of $F_{m}(k>k_{c})=0$ with localized systems for a certain finite $k_{c}$ value. Conversely, we associate $F_{m}(k)>0$ with delocalized systems.
@@
In Fig. 1 (d) we examine $F_{m}(k)$ as $k$ increases. When $\Gamma>0$ and diffusion occurs, $F_{m}(k)$ converges to a positive number as $k\to\infty$ which, although not directly proportional to it, correlates with the non-monotonic behavior of the diffusion constant with $\Gamma$, behavior which is well understood.[24, 54, 55] Both the diffusion constant and $F_{m}(k)$ share a similar maximum location near $\Gamma=1$ for the chosen disorder strength. Note that the outgoing-pseudoflux $f_{m}(k)$ is always positive when $k=1$ and negative otherwise, indicating that the role of memory is to return the population back to its source, suppressing diffusion. This is not surprising given that removing memory is equivalent to introducing $\Gamma$ and destroying localization. We note that a different and perhaps more seamless connection between disorder and noise can be achieved with spatially uncorrelated but temporally correlated noise. Static disorder in this sense can be thought of as perfectly and indefinitely correlated noise while white noise is the opposite limit. In Appendix B we examine systems with exponentially correlated noise $\langle\epsilon_{n}(t)\epsilon_{m}(t^{\prime})\rangle=\Gamma e^{-\gamma t}\delta_{nm}$, in which the two limits can be continuously bridged with a single parameter $\gamma$, the inverse correlation time. Similar conclusions on memory and transport can be drawn there.
@@
Figure 2: All calculations in this figure were run with $\Delta t=V/2$ and 20002000 trajectories and, unless otherwise specified, a basis of size 2121. (a) The Frobenius norm of the transfer tensor $T(k)||$ as a function of time step $k$ for several different $\lambda$ strengths. Note that the norm of $T(1)$ is on the order of the basis size and not shown on these axes. (b) The cumulative outgoing-psuedoflux after $500$ time steps as a function of $\lambda$ for several different basis sizes of the AAH model, showing convergence towards the expected behavior of an infinite size system as the basis size increases. The inserted plot shows the convergence towards the final cumulative values as a function of time step for the varying basis sizes at a fixed $\lambda=1.5$. Note that the cumulative outgoing-pseudoflux for $T(1)$ is not shown on these axes. (c) The Frobenius norm of the transfer tensor as a function of index given different $\Gamma$ for an AAH model with $\lambda=2$, with $T(1)$ again not shown on these axes. (d) The cumulative outgoing-pseudoflux as a function of time step for the same models as in (c).
@@ III Transfer Tensor Analysis of Aubry-André-Harper Model
To test this, we turn our attention to the Aubry-André-Harper (AAH) model, [17, 18] which has been physically simulated with optical lattices.[38, 39, 40, 41, 42, 43] In the AAH model, instead of a random distribution, the site energies are given by $\epsilon_{n}=\lambda\cos(2\pi\beta n+\phi)$, where $\lambda$ is the strength of the potential, $\beta$ is an irrational number, the golden ratio in our case, and $\phi$ is a phase factor. Here, the averaging is performed with respect to $\phi$, which is drawn from a uniform distribution spanning $0,2\pi]$. The eigenstates are localized above a critical value of $\lambda=2V$ and delocalized below it.[29, 17] We immediately see that, as shown in Fig. 2 (a), this model displays infinite memory time regardless of $\lambda$. Infinite memory time does not provide an indication of localization, although a larger $\lambda$ results in convergence of the transfer tensor norm to a larger value.
At approximately $\lambda/V=2$ in Fig. 2 (b), there is an obvious change in the behavior of $F_{m}(k)$, although the basis size has an influence on the localization to delocalization behavior. The cumulative outgoing-pseudoflux at the end of the simulation is shown as a function of $\lambda$ for several basis sizes. Smaller bases show less obvious delocalization to localization transitions, and these transitions appear at smaller values of $\lambda$. This is likely a result of density returning to its origin after reflecting off the system boundary. This leads to an underestimate of $\lambda$ for the localization transition in finite systems. Similarly, care must be taken in the Anderson model when $\sigma$ is so small that the eigenstate localization length exceeds the size of the finite basis. The largest basis size employed for the AAH model, $21$ in Fig. 2 (b) is sufficiently converged towards the infinite limit to clearly indicate the transition point at $\lambda/V=2$. An inset in Fig. 2 (b) shows how $F_{m}(k)$ converges to finite values as a function of $k$ for the case of $\lambda/V=1.5$.
HSR white noise can be introduced to the AAH model in precisely the same manner as in the Anderson model.[59] Inclusion of heat baths in this model has been previously reported to induce transport.[60] Introduction of noise results in a finite memory time, as demonstrated in Fig. 2 (c). The AAH model also shows the same kind of non-monotonic behavior of $F_{m}(k)$ with increasing $\Gamma$ as seen in the Anderson model, with the outgoing-pseudoflux converging to non-zero values which are maximized at a moderate value of $\Gamma\approx 1$, suggesting non-monotonic behavior of transport efficiency with noise.

### v2

@@ Abstract
We use the transfer tensor method to analyze localization and transport in simple disordered systems, specifically the Anderson and Aubry-André-Harper models. Emphasis is placed on the memory effects that emerge when ensemble-averaging over disorder, even when individual trajectories are strictly Markovian. We find that transfer tensor memory effects arise to remove fictitious terms that would correspond to redrawing static disorder at each time step, which would create a temporally uncorrelated dynamic disorder. Our results show that while eternal memory is a necessary condition for localization, it is not sufficient. We determine that signatures of localization and transport can be found within the transfer tensors themselves by defining a metric called “outgoing-pseudoflux”. This work establishes connections between theoretical research on dynamical maps and Markovianity and localization phenomena in physically realizable model systems.
@@ I Introduction
Theoretical study of quantum transport in realistic systems is complicated by the unfavorable scaling of computational requirements when modeling quantum systems,[15] so simple but powerful treatments such as the Anderson,[3], Haken-Strobl-Reineker (HSR),[59, 35] or Aubry-André-Harper (AAH) models[4, 36] are often employed to understand transport and localization in tight binding lattices. As many experimental measurements sample over regions with defects, inducing effects such as inhomogeneous broadening on spectroscopic measurements,[38, 10, 56, 7, 64] it is common practice to average over static disorder in model parameters in order to better describe experimental observations.[3, 46, 52, 31] These simple models have increased our understanding of phenomena including metal-insulator transitions and diffusion rates in the presence of thermal noise and disorder.[53, 68, 67, 46, 28, 4, 36, 52]
The dynamics resulting from averaging over a static disorder ensemble recovers the spatial homogeneity that was destroyed when the disorder was introduced. However, a somewhat counterintuitive complication arises. Evolutions consisting of mixtures of distinct unitary evolutions, averages over static disorder in our case, may show memory effects. Although the exact interpretation of these effects, and of quantum non-Markovianity in general, remains a matter of discussion,[13, 37, 17, 5, 50] the association of this apparent non-Markovianity and Anderson localization has been noted in previous studies.[48, 47, 40] We will reserve the connection of this phenomenon to other established measures of quantum non-Markovianity and its potential connection to information backflow for future studies,[61, 14, 13] and rather analyze the origin of these memory effects using transfer tensors (TT), which are ideally adapted to this purpose. Transfer tensors or related process tensors have proved useful characterizing environmental noise,[55, 22, 24, 23] extracting information about emitted photons,[26] and making arguments about the divisibility of dynamical maps characterizing system evolution.[32] Here we restrict ourselves to stating that a system has memory when memory terms appear in the transfer tensor hierarchy without any intended implication about non-Markovianity, such as that arises from genuine non-perturbative system-bath interactions.  [^1] We note that as discussed in Ref. [32], having vanishing $k>1$-step tensors is a sufficient condition for CP-divisible evolutions and the dynamics is consequently Markovian by that criterion. Here, the TTs from mixtures of Markovian dynamics are (uniquely) constructed identically to that used in Ref. [32]. However, there is no system-environment correlation in the traditional open quantum system sense We see that memory terms differentiate static disorder from dynamic disorder and find that, although the existence of memory does not necessarily lead to localization, another characteristic of the transfer tensors, which we call outgoing-pseudoflux, provides an indication of localization. This demonstrates the potential of TT as a diagnostic tool for assessing transport properties and understanding the origin and impact of memory derived from disorder averaging.
@@ II Transfer Tensor Analysis of the Anderson Model
in the lattice site basis, $|n\rangle$, which includes nearest-neighbor interactions with coupling strength $V$ and site energies, $\epsilon_{n}$, drawn from a normal distribution with zero mean and standard deviation $\sigma$. Throughout this paper, we set $\hbar=1$ and $V=1$. Introducing static disorder localizes the system in 1D and breaks spatial homogeneity,[3, 58] but averaging over many realizations of the static disorder restores it. Whereas the individual trajectories in the ensemble are Markovian and have no memory, the spatially homogeneous ensemble evolution now has memory, as noted in the context of convex mixing of quantum channels or dynamical maps[15, 13, 37] and in previous studies of Anderson localization.[48, 47] To better see this, we adopt the transfer tensor method to describe the disorder-averaged dynamics.
@@
We will find that individual elements of the transfer tensor can provide useful information. Because we average over disorder and restore spatial homogeneity, we can analyze only the transfer tensor elements involving the center site of the system. The element $T_{m,n,o,p}(l)$ quantifies the influence that $\rho_{o,p}(k)$ will have on $\rho_{m,n}(k+l)$ and $T_{m,n,o,p}(l)$ depends only on the differences mpm-p, npn-p, and opo-p. The utility of an equivalent relationship for the generalized quantum master equations was recently recognized and exploited for performance enhancement. [9] The spatial homogeneity of the TT allows analysis with modest basis sizes.
@@
Figure 1: (a) The Frobenius norm of the transfer tensor $T(k)||$ as a function of $k$ for $\Gamma=0$ with different $\sigma$ values. (b) Population of site $m=8$ following initialization in a delta function localized on $m$ as a function of time step for several ensembles including a case in which all transfer tensor memory terms were artificially set to zero. The dashed black line is the equilibrium population. (c) The Frobenius norm of the transfer tensor $T(k)||$ as a function of time step $k$ for several different $\Gamma$ values with $\sigma=V$. (d) The cumulative outgoing-pseudoflux as a function of time step for the same $\Gamma$ and $\sigma$ values as in (c) showing convergence to non-zero values at finite $\Gamma$ and convergence to zero at $\Gamma=0$. Note that, when $\Gamma=10V$, noise overpowers disorder and the system is effectively memoryless for the chosen $\Delta t$. All calculations in this figure involved a basis of size 1515, $V\Delta t=0.1$ and 20002000 trajectories. In (a) and (c) the norm of the first transfer tensor is on the order of the basis size and not shown on these axes.
The memory that arises for the disorder-averaged ensembles is quantified in Fig. 1 (a) which shows the Frobenius norms of the TT, $T(k)||$, as a function of $k$ for several $\sigma$. In the case where $\sigma=0$, memory is zero within the numerical tolerance for all $k>1$. For $\sigma>0$, the tensor norms converge toward a finite value as $k\to\infty$. This finite value increases with $\sigma$ and suggests an eternal memory, in the sense that there will be an indefinite number of memory terms in the transfer tensor. A similar phenomenon has been noted in the study of the eternally non-Markovian quantum master equation, which is characterized by a quasi-Gorini-Kossakowski-Sudarshan-Lindblad[45, 33] form in which there is a perpetually negative rate associated with a jump operator, which is taken as an indication of quantum non-Markovianity.[50] It is important, however, to point out that the ensemble evolutions we address can be viewed as convex mixtures of Markovian maps. There are competing definitions of quantum non-Markovianity and there has been debate as to whether convex combinations of Markovian maps should be considered to demonstrate it.[39, 50, 17, 13, 43, 37, 40, 18] Specifically, information backflow as a measure of non-Markovianity has been debated in this context.[50, 13, 5] Whether the ensemble averaged systems we analyze should be considered to display quantum non-Markovianity or whether the appearance of memory terms in the transfer tensor in general is indicative of quantum non-Markovianity is beyond the current scope and we will avoid using this term in favor of discussing memory. When we refer to memory in this work, we refer only to the presence of transfer tensor memory terms, that the Frobenius norm of any $k>1$-step TT is finite.
@@
Upon closer examination, the Markovian evolution represented by $M(k)=[T(1)]^{k}$ in the above discussion is equivalent to the dynamics induced by averaging uncorrelated Gaussian white noise in the limit of $\Delta t\rightarrow 0$. This is actually the HSR model, averaged over stochastic noise,[35, 59] where the site energies follow zero mean Gaussian statistics with no spatial or temporal correlations
@@
This gives rise to the equivalent Gorini–Kossakowski–Sudarshan–Lindblad equation of motion for the corresponding noise-averaged density matrix,[45, 33]
@@
When the HSR model is combined with the disordered Anderson Hamiltonian the dynamics shows diffusive transport whenever $\Gamma>0$.[52, 49, 57, 41] We examine the interplay between static disorder and dynamic noise with respect to transport further using the TT formalism. One can readily deduce that in the limit of an infinitesimal time step, $M(1)$ will be identical for all ensembles for which $\Gamma+\sigma^{2}\Delta t=CV$, with $C$ an arbitrary positive constant. Figure 1 (b) shows the population as a function of time at the center site for a set of ensembles with $C=0.1$, in which all dynamics are initialized as a Kronecker delta. In one special case, an ensemble was prepared for $\sigma=V$ and $\Gamma=0$, but the resulting memory terms were artificially set to zero, i.e. $T(k)=0$ for all $k>1$ while keeping the first term the same. The resulting dynamics is identical to a $\sigma=0$, disorder-free and naturally memoryless propagation with $\Gamma=0.1V$. This demonstrates that when memory terms are eliminated, static and dynamic disorder cannot be differentiated. In Fig. 1 (b), several time steps are required before memory effects begin to manifest impacts on M(i)M{(i)} and the populations from different ensembles diverge, with divergence occurring more quickly for larger $\sigma$. Regardless of the size of $\sigma$, as long as $\Gamma$ is nonzero all populations will eventually converge to the infinite temperature equilibrium state where all sites are equally occupied, indicated with a dashed black line in the figure. In the case in which $\sigma=V$ and there is no noise, the population on the site where the system was initialized remains above the equilibrium value.
@@
Second, inspection of individual tensor elements can indicate transport properties. It is customary to follow the time-dependent spatial profile with certain localized initial conditions as in microscopy experiments to assess transport.[65, 51, 10] To facilitate comparisons between different systems, averaged measures such as mean squared displacement $\langle x^{2}\rangle$, whose scaling with time can be used to classify the nature of transport, e.g. $\langle x^{2}\rangle\propto t$ for diffusive transport, are employed. Here, we take a different approach and define the outgoing-pseudoflux at time step $k$ for site $m$ as $f_{m}(k)=\sum_{n}^{\prime}{T}_{n,n,m,m}(k),\;\;n\neq m$ and its cumulative sum over time $F_{m}(k)=\sum_{l=1}^{k}f_{m}(l)$. The former represents the magnitude of the influence that the population of the $m$th site $k$ steps in the past has on the populations of all other sites in the system at the current time step. The physical significance of $F_{m}(k)$ is further discussed in Appendix A, where we illustrate the association of $F_{m}(k>k_{c})=0$ with localized systems for a certain finite $k_{c}$ value. Conversely, we associate $F_{m}(k)>0$ with delocalized systems.
@@
In Fig. 1 (d) we examine $F_{m}(k)$ as $k$ increases. When $\Gamma>0$ and diffusion occurs, $F_{m}(k)$ converges to a positive number as $k\to\infty$ which, although not directly proportional to it, correlates with the non-monotonic behavior of the diffusion constant with $\Gamma$, behavior which is well understood.[52, 60, 19] Both the diffusion constant and $F_{m}(k)$ share a similar maximum location near $\Gamma=V$ for the chosen disorder strength. Note that, in the current model, the outgoing-pseudoflux $f_{m}(k)$ is always positive when $k=1$ and negative otherwise, indicating that the role of memory is to return the population back to its source, suppressing diffusion. This is not surprising given that removing memory is equivalent to introducing dynamical noise and destroying localization. We note that a different and perhaps more seamless connection between disorder and noise can be achieved with spatially uncorrelated but temporally correlated noise. Static disorder in this sense can be thought of as perfectly and indefinitely correlated noise while white noise is the opposite limit. In Appendix B we examine systems with exponentially correlated noise $\langle\epsilon_{n}(t)\epsilon_{m}(t^{\prime})\rangle=\Gamma e^{-\gamma t}\delta_{nm}$, in which the two limits can be continuously bridged with a single parameter $\gamma$, the inverse correlation time. Similar conclusions on memory and transport can be drawn there.
@@
Figure 2: (a) The Frobenius norm of the transfer tensor $T(k)||$ as a function of time step $k$ for several different $\lambda$ strengths. (b) The cumulative outgoing-psuedoflux after $500$ time steps as a function of $\lambda$ for several different basis sizes of the AAH model, showing convergence towards the long-chain limit. The inset shows the convergence towards the final cumulative values as a function of time step for the varying basis sizes at a fixed $\lambda=1.5V$. (c) The Frobenius norm of the transfer tensor as a function of index given different $\Gamma$ for an AAH model with $\lambda=2V$. (d) The cumulative outgoing-pseudoflux as a function of time step for the same models as in (c).All calculations in this figure were run with $V\Delta t=1/2$ and 20002000 trajectories and, unless otherwise specified, a basis of size 2121. Note that the norm of $T(1)$ is on the order of the basis size and not shown on the axes of (a) and (c).
@@ III Transfer Tensor Analysis of Aubry-André-Harper Model
To test this, we turn our attention to the Aubry-André-Harper (AAH) model, [4, 36] which has been physically simulated with optical lattices.[6, 62, 42, 27, 11, 44] In the AAH model, instead of a random distribution, the site energies are given by $\epsilon_{n}=\lambda\cos(2\pi\beta n+\phi)$, where $\lambda$ is the strength of the potential, $\beta$ is an irrational number, the golden ratio in our case, $\phi$ is a phase factor, and V-V replaces $V$ in Eq. 1. Here, the averaging is performed with respect to $\phi$, which is drawn from a uniform distribution spanning $0,2\pi]$. The eigenstates are localized above a critical value of $\lambda=2V$ and delocalized below it.[28, 4] We immediately see that, as shown in Fig. 2 (a), this model displays eternal memory regardless of $\lambda$. Such memory does not provide an indication of localization, although a larger $\lambda$ results in convergence of the transfer tensor norm to a larger value.
At approximately $\lambda=2V$ in Fig. 2 (b), there is a notable change in the behavior of $F_{m}(k)$, although the basis size has an influence on the localization to delocalization behavior. The cumulative outgoing-pseudoflux at the end of the simulation is shown as a function of $\lambda$ for several basis sizes. Smaller bases show less obvious delocalization to localization transitions, and these transitions appear at $\lambda<2V$. This is likely a result of density returning to its origin after reflecting off the system boundary. This leads to an underestimate of $\lambda$ for the localization transition in finite systems. Similarly, care must be taken in the Anderson model when $\sigma$ is so small that the eigenstate localization length exceeds the size of the finite basis. The largest basis size employed for the AAH model, $21$ in Fig. 2 (b) is sufficiently converged towards the long-chain limit to clearly indicate the transition point at $\lambda=2V$. An inset in Fig. 2 (b) shows how $F_{m}(k)$ converges to finite values as a function of $k$ for the case of $\lambda=1.5V$.
HSR white noise can be introduced to the AAH model in precisely the same manner as in the Anderson model.[noteone] Inclusion of heat baths in this model has been previously reported to induce transport.[8] Introduction of noise results in a finite memory time, as demonstrated in Fig. 2 (c). The AAH model also shows the same kind of non-monotonic behavior of $F_{m}(k)$ with increasing $\Gamma$ as seen in the Anderson model, with the outgoing-pseudoflux converging to non-zero values which are maximized at a moderate value of $\Gamma\approx V$, suggesting non-monotonic behavior of transport efficiency with noise.